OpenStudy (anonymous):

For the function shown below, find f(2): 2x4-8x2+5x-7

OpenStudy (anonymous):

That's math

OpenStudy (anonymous):

F(2)=2(2)^4-8(2)^2+5(2)-7 =32-32+10-7 =3

OpenStudy (anonymous):

Thanks Got it right!